Suppose demand and supply are given by Qd = 60 – P and Qs = P...

  1. Suppose demand and supply are given by Qd = 60 – P and Qs = P -20
  1. What are the equilibrium quantity and price in this market?
  2. Determine the quantity demanded, the quantity suppled, and the magnitude of the surplus id a price floor of $50 is imposed in this market.
  3. Determine the quantity demanded, the quantity suppled, and the magnitude of the shortage if a price celling of $32 is imposed in this market. Also determine the full economic price paid by consumers.
Answer #1

a. The equilibrium is where the quantity demanded equals the quantity supplied. here, the equilibrium thus is where 60-P=P-20

2P=80

P=$40

The corresponding equilibrium quantity is 60-40=20 units

b. when there is a price floor of $50 being imposed, the total quantity demanded then will be 60-50=10 while the total quantity supplied will be 50-20=30. Since the supply is greater than the demand, thus we can say that there is a surplus in the market by 30-10= 20 units.

c. Similarly, when the price ceiling of $32 is imposed in the market, the total quantity demanded then will be 60-32=28 while the total quantity supplied will be 32-20=12. Since the supply is lesser than the demand, thus we can say that there is a shortage in the market by 28-12= 16 units

The full economic price is the price where the demand meets the supply at the shortage amount. here since the shortage is of 16 units, thus the full economic price paid by the consumers is Qd=16=60-P

P=60-16=$44